# FiDeLiS: Faithful Reasoning in Large Language Models for Knowledge Graph Question Answering

### v2

@@ 3 Experiments
Prompting - LLM Only gpt-4-turbo Zero-shot 62.32 59.71 42.71 37.93 37.74
@@
Finetuning - LLM + KG NSM (He et al., 2021) 74.31 - 53.92 - -
CBR-KBQA (Das et al., 2021) - - 67.14 - -
DeCAF (Yu et al., 2022) 82.1 - 70.42 - -
KD-CoT (Wang et al., 2023a) 73.7 50.2 50.5 - -
RoG (Luo et al., 2024) 83.15 69.81 61.39 56.17 60.32
Prompting - LLM + KG gpt-3.5-turbo ToG (Sun et al., 2023) 75.13 72.32 57.59 56.96 62.48
@@
3.1 RQ1: KGQA Performance Comparison
Table 1 clearly showcases the comparative effectiveness of FiDeLiS, against different baselines. It shows that FiDeLiS outperforms all baselines using gpt-4-turbo model, even for strong baselines being fine-tuned, such as DeCAF (Yu et al., 2022) and RoG (Luo et al., 2024). Across all prompting-based methods, gpt-4-turbo achieves higher performance compared to gpt-3.5-turbo, especially on the CWQ dataset. It indicates that gpt-4-turbo has a better understanding and processing capability for complex queries. The CR-LT dataset seems more challenging compared with WebQSP and CWQ, as implied by the consistently lower accuracy scores. However, FiDeLiS shows consistent enhancements compared to other baselines, demonstrating its capability to handle long-tail entities by referring to the knowledge from KGs and processing more complex queries effectively.
3.2 RQ2: Ablation Study of FiDeLiS.
Table 2 demonstrates the ablation study of FiDeLiS using the gpt-3.5-turbo-0125 model across the WebQSP, CWQ, and CR-LT datasets. It highlights the critical importance of both the Path-RAG and DVBS components. Removing Path-RAG, whether by employing a vanilla retriever or the think-on-graph (ToG) approach (Sun et al., 2023), results in substantial performance declines, particularly a 6.97% drop in Hits@1 on WebQSP and a 6.01% decrease on CWQ, underscoring the necessity of an effective retrieval mechanism. Similarly, ablating parts of DVBS, especially the beam search component, leads to significant reductions, with an 18.97% decrease in Hits@1 on WebQSP and a 13.34% drop on CWQ, indicating that beam search is vital for maintaining high accuracy. Other DVBS subcomponents, such as the deductive verifier and planning, also contribute notably to performance, albeit to a lesser degree. Overall, the findings demonstrate that the integrated Path-RAG and DVBS frameworks are essential for FiDeLiS’s robust performance, with their removal causing marked decreases in accuracy and Hits@1 scores across all evaluated datasets.
Table 2: Ablation Studies of FiDeLiS using model gpt-3.5-turbo-0125. $\Delta$ refers to the performance gap between each component and the entire method. The three largest performance gaps on each dataset are highlighted in green, with darker shades denoting more significant differences.
@@ 3.1 RQ1: KGQA Performance Comparison
Hits@1 (%) $\Delta$ Hits@1 (%) $\Delta$ Acc (%) $\Delta$
No ablation FiDeLiS 79.32 0.00 63.12 0.00 67.34 0.00
w/o Path-RAG using vanilla retriever 72.35 6.97 57.11 6.01 59.78 7.56
using ToG 75.11 4.21 59.47 3.65 63.47 4.07
w/o DVBS w/o last step reasoning 75.68 3.64 59.45 3.67 63.72 3.62
w/o planning 76.23 3.09 60.14 2.98 64.13 3.21
w/o beam-search 60.35 18.97 49.78 13.34 61.87 5.47
w/o deductive-verifier 74.13 5.19 57.23 5.89 63.89 3.45
3.3 RQ3: Robustness Analysis
Comparison using different embedding backbones. Table 3 demonstrates the performance of the Path-RAG using various embedding models as backbones. It shows that the integration of the Openai-Embedding Model significantly enhances the performance across all datasets, indicating its superior capability in understanding and processing natural language queries, which leads to more effective retrieval. For instance, the Openai-Embedding Model increases the Hits@1 by 13.04% in the WebQSP dataset and 14.73% in the CWQ dataset compared to the base BM25 (Robertson et al., 2009) model. Furthermore, this improvement is consistent and highlights the feasibility and potential of pairing more powerful language models with structured reasoning frameworks to address complex information retrieval and question answering challenges.
Table 3: Performance FiDeLiS using different Path-RAG embedding backbone models. ‘Openai-Embedding Model’ refers to text-embedding-3-small, the recent embedding model released from OpenAI.
@@ 3.3 RQ3: Robustness Analysis
Vanilla Retriever w/ BM25 (Robertson et al., 2009) 58.31 48.39 50.73
w/ SentenceBert (Reimers & Gurevych, 2019) 62.74 50.14 51.80
w/ E5 (Wang et al., 2022a) 68.42 52.84 54.31
w/ Openai-Embedding-Model 72.35 57.11 59.78
Path-RAG w/ BM25 (Robertson et al., 2009) 70.34 56.11 58.77
w/ SentenceBert (Reimers & Gurevych, 2019) 73.45 58.41 60.45
w/ E5 (Wang et al., 2022a) 77.93 62.74 65.23
w/ Openai-Embedding-Model 79.32 63.12 67.34
Comparison using different searching widths and depths. To investigate how the search width and depth affect the performance of FiDeLiS, we have performed experiments with settings ranging from depths of 1 to 4 and widths of 1 to 4. The results, presented in Figure 2, show that performance generally improves as search depth and width increase. However, performance begins to decline when the search depth exceeds 3 for both the WebQSP and CWQ datasets. This decline is attributed to the fact that only a small fraction of questions in these datasets require reasoning at depths greater than 3. Additionally, FiDeLiS utilizes stricter rules for ending the search process, which can preemptively terminate misguided searches. Contrastingly, increasing the search width consistently shows potential for improved performance, suggesting benefits from broader exploration. However, considering the computational costs, which rise linearly with depth, we have chosen to set both the default beam width and depth to 4 for all experiments unless stated otherwise. This default setting aims to balance performance gains with computational efficiency.
(a) RD over CWQ
(b) RD over WebQSP
(c) BW over CWQ
(d) BW over WebQSP
Figure 2: Performance (Hits@1) of FiDeLiS with different beam search widths (BW) and reasoning depths (RD) over WebQSP and CWQ.
Comparison between Path-RAG and Vanilla Retriever. To verify the robustness of the Path-RAG compared to vanilla retrievers. We conduct the experiments as shown in Figure 3(a) and 3(b). We calculate the coverage ratio (CR) of the retrieved reasoning paths and the ground-truth reasoning paths as follows: $\text{CR}=\left(\frac{N_{\text{retrieved}}\cap N_{\text{ground-truth}}}{N_{\text{ground-truth}}}\right)\times 100\%$, where $N_{\text{retrieved}}$ is the number of retrieved reasoning paths candidates, and $N_{\text{ground-truth}}$ is the total number of ground-truth reasoning paths. We set the vanilla retriever as the baseline, specifically, we concatenate each entity with its corresponding relation to form a reasoning step. We then calculated the cosine similarity between the embeddings of each reasoning steps and the query embeddings to select candidate paths. We observe that compared to the vanilla retriever, Path-RAG achieves a higher CR value and shows better alignment with the ground-truth paths. It demonstrates that our method can better leverage the connections that are often missed by simpler retrieval models. This feature is critical in subsequent LLM to focus on the information of interest and arrive at a correct and reasonable answer.
(a) Coverage Ratio over CWQ
(b) Coverage Ratio over WebQSP
(c) Error Analysis over WebQSP
Figure 3: (a)-(b): Empirical study regarding the coverage ratio of the retrieved reasoning paths. (c): Error analysis of the validation of whole paths generated from RoG (Luo et al., 2024) over WebQSP.
@@
Error analysis regarding whole path generation. To verify the faithfulness of our stepwise method, we conduct the error analysis regarding the whole reasoning path generation. Specifically, we conduct an analysis of the validity of the whole reasoning path generation using the baseline methods RoG (Luo et al., 2024). The result is shown in Figure 3(c). The definition of the metric validity ratio (VR) is the ratio of reasoning steps that existed in the knowledge graph to the total number of the reasoning steps in the output reasoning path: $\text{VR}=\left(\frac{N_{\text{valid-steps}}}{N_{\text{all-steps}}}\right)\times 100\%$, where $N_{\text{valid-steps}}$ is the number of reasoning steps that existed in the knowledge graph, and $N_{\text{all-steps}}$ is the number of all the reasoning steps in the reasoning path. It show that only 67% of generated reasoning steps are valid, while the remaining 33% of reasoning steps either have a format error or do not exist in the KG. This illustrates that the reasoning steps generated by language models offer few guarantees about feasibility especially when multiple consecutive steps are combined into a reasoning path.
Table 4: Average depths of the returned reasoning paths. GT refers to ground-truth reasoning paths.
@@
Effectiveness of deductive-verification. To verify the effectiveness of deductive-verification, we calculate the average depths of the grounded reasoning path from ToG and our methods in Table 4. It shows that FiDeLiS consistently shows shorter and closer reasoning depths to ground-truth across all datasets compared to ToG. This implies that our method may offer more precise termination signals and potentially more accurate reasoning paths compared to baselines.
Case Study. To verify whether FiDeLiS can conduct faithful reasoning on real cases, we also present a case study in Table 5. We can find that FiDeLiS illustrates better performance in identifying multiple possible governmental forms—Theocracy, Unitary state, and Islamic republic established in Iran in 1979, as compared to existing models such as GPT + CoT, RoG, and ToG. This observation highlights the importance of recognizing the complexity of Iran’s government system, which combines elements of religion and democracy. Different models handle this complexity with varying levels of detail. For example, GPT + CoT and RoG mainly identified Iran as an ‘Islamic Republic’, a change that occurred after the 1979 Iranian Revolution, but they missed other possible descriptions. The ToG model even gives mixed results, such as simply naming it ‘Iran’, which shows confusion in its analysis. However, our method shows a better understanding of the context and provides a more detailed and accurate set of answers. This illustrates FiDeLiS strength and efficiency in dealing with intricate questions.
Table 5: Case study of FiDeLiS. The wrong answers are highlighted with red color, and the correct answers are highlighted with blue color. In this case, FiDeLiS scored an F1 of 0.857, outperforming the RoG method (0.4).
3.4 RQ4: Efficiency of FiDeLiS
To investigate the runtime efficiency and cost efficiency of FiDeLiS, we present a comparison regarding the average runtime, average token usage, average times of LLM calling per question in Table 6. We found that (1) our method shows superior efficiency compared to the ToG (which is also training-free), by reducing approximately 1.7x runtime costs. (2) Path-Rag component is critical in enhancing both the accuracy and efficiency of the model. Its ability to constrain potential path candidates effectively reduces unnecessary computational overhead, leading to quicker and more accurate results.
Table 6: Runtime efficiency of FiDeLiS per question.
Dataset Method Hits@1 (%) Avg Runtime (s) Avg Token Usage Avg LLM calling
@@ 3.4 RQ4: Efficiency of FiDeLiS
To address concern regarding our method’s potential application in real-time scenarios, we also tested our method using faster and more advanced LLMs. Table 6 shows that our method could be further accelerated with newer, faster models like GPT-4o or GPT-4-mini. The potential of the ongoing advancements in LLMs are expected to further enhance the scalability and efficiency of FiDeLiS, making it a practical development in challenging environments. More detailed analysis of bottleneck of computation of FiDeLiS can be further found in Appendix C.
@@ 4 Related Work
KG-enhanced LLM. KGs have advantages in dynamic, explicit, and structured knowledge representation and techniques combining LLMs with KGs have been studied (Pan et al., 2024). Early studies (Luo et al., 2024; Yu et al., 2022) embed structured knowledge from KGs into the underlying neural networks during the pretraining or fine-tuning process. However, the reasoning steps generated by LMs are observed to be prone to errors, which could be non-existent in the knowledge graph, and can subsequently lead to incorrect reasoning during inference. Also, KG embedded in LLM sacrifices its own nature of explainability in knowledge reasoning and efficiency in knowledge updating (Hu et al., 2023). Another solutions are to keep the reasoning model as an agent to explore the external structure knowledge source (Sun et al., 2023; Gu et al., 2023; Jiang et al., 2023b; Wang et al., 2023a) ToG (Sun et al., 2023) directly employs LLMs to output scores for candidate selections from the KG. This model operates at the relation-entity level of the KG, aiming to identify relevant triples that aid the LLM in making accurate and responsible final answer predictions.
Knowledge Graph-based Question Answering. To integrate LLMs for KGQA, retrieval-augmented methods aim to retrieve the relative facts from the KGs to improve the reasoning performance (Li et al., 2023; Karpukhin et al., 2020). Recently, UniKGQA (Jiang et al., 2022) which unifies the graph retrieval and reasoning process into a single model with LLMs, achieves STOA performance. Semantic parsing methods convert the question into a structural query (e.g., SPARQL) by LLMs, which can be executed by a query engine to reason the answers on KGs (Sun et al., 2020; Lan & Jiang, 2020). However, these methods heavily rely on the quality of generated queries. If the query is not executable, no answers will be generated. DECAF (Yu et al., 2022) combines semantic parsing and LLMs reasoning to jointly generate answers, which also reach salient performance.

### v3

@@ 3.2 Deductive-Verification Beam Search
[…]
Prompting - LLM Only gpt-4-turbo IO 62.32 59.71 42.71 37.93 37.74
@@
[…]
Finetuning - LLM + KG NSM He et al. (2021) 74.31 - 53.92 - -
CBR-KBQA Das et al. (2021) - - 67.14 - -
DeCAF Yu et al. (2023) 82.1 - 70.42 - -
KD-CoT Wang et al. (2023) 73.7 50.2 50.5 - -
RoG Luo et al. (2024a) 83.15 69.81 61.39 56.17 60.32
[…]
Prompting - LLM + KG gpt-3.5-turbo ToG Sun et al. (2023) 75.13 72.32 57.59 56.96 62.48
KAPING Baek et al. (2023) 72.42 65.12 53.42 50.32 -
@@
Beam Search.
We then construct the multi-step reasoning paths by iteratively extending partial paths using a beam search strategy. At each time step $t$, we use an LLM as agent to select one reasoning step $s^{t}$ from a candidates set $\mathcal{S}^{t}$3.1) conditioned on (1) the likelihood of each candidate $s_{i}\in\mathcal{S}^{t}$, (2) the user query $q$, (3) the history of previous steps $s^{1:{t-1}}$, and (4) planning context $w$3.2), denoted as $\mathrm{LM}(s^{t}|q,s^{1:t-1},w,\mathcal{S}^{t})$. Instead of exploring every possibility, we retain only the top-$k$ scoring paths from the previous beam $\mathcal{H}_{t-1}$ and extend them by appending candidate steps. The overall process can be expressed as Eq 3:
$\mathcal{H}_{t}=\mathrm{Top}_{k}\big{\{}h\oplus\mathrm{LM}(s^{t}|q,s^{1:t-1},w,\mathcal{S}^{t}):h\in\mathcal{H}_{t-1}\big{\}}$ (3)
where $\oplus$ denotes the concatenation of the current path $h$ with the selected candidate step $s^{t}$. The beam search strategy enable efficiently navigate the vast space of potential reasoning paths while concentrating on the most promising ones.
While beam search, by its nature, can incur high computational costs and latency due to multiple rounds of LLM interactions. Our retrieval module Path-RAG mitigate this issue by constraining candidate set $\mathcal{S}^{t}$ at each time step $t$ to a narrow, high-quality subset rather than requiring the LLM to consider all available options. This targeted retrieval not only reduces the number of candidates to evaluate at each step but also increases the likelihood of selecting relevant reasoning steps, thereby enabling efficient traversal of KGs at scale. Find more discussion regarding efficiency of FiDeLiS in §4.4 and Appendix §B.
Deductive Verification.
To ensure that each reasoning step logically follows from its predecessors and adequately supports the original query, we leverage the deductive reasoning capabilities of LLMs as a verification criterion Ling et al. (2023) for the beam search process. We first convert the user query $q$ into a clear declarative statement $q^{\prime}$, which encapsulates its logical intent and allows the LLM to operate on a well-defined logical target (See the concrete example in §C.6). Next, during the beam search, candidate reasoning step $s^{t}$ are appended to the history $s^{1:t-1}$ to form potential reasoning paths. For each candidate, we then invoke two deductive verification checks, $C_{\mathrm{global}}$ and $C_{\mathrm{local}}$ (the prompts are given in §C.2). Only those candidates that pass local verification, indicating that the new step maintains logical consistency with the established context, are retained in the beam search process. Once the candidates pass both verification indicate that the user query $q$ can be deduced based on the retained reasoning paths $s^{1:t}$ and the beam search progress should be halted.
Global Verification: $C_{\mathrm{global}}(s^{1:t-1},s^{t})$ returns 1 if $(s^{t}\land s^{1:t-1})\models q^{\prime}$, and 0 otherwise.
Local Verification: $C_{\mathrm{local}}(s^{1:t-1},s^{t})$ returns 1 if $s^{t}$ logically follows from $s^{1:t-1}$, and 0 otherwise.
By integrating this verification into the beam search offers several benefits: it (1) enhances the robustness and validity of the final answer by enforcing logical coherence at every step, (2) reduces computational overhead by pruning unpromising paths early, and (3) mitigates risks such as premature termination or excessive extension of the reasoning process. We provide a concrete example of the deductive verification process in §C.6 and the complete DVBS algorithm in Algorithm 2.
4 Experiments
In this section, we focus on verifying FiDeLiS from four perspectives as follows: (1) comparison results with other baselines over KGQA; (2) ablation study; (3) robustness analysis and (4) efficiency analysis. We provide all the experiment settings in Appendix A due to page constraints. The prompts for plan generation, beam search and deductive verification can be found in §C.
4.1 Main Results
In Table 1, we compare the performance of different methods with various backbend LLMs across three datasets. We found that LLM + KG approaches generally outperform LLM-only methods (Zero-shot, Few-shot, and CoT) by a wide margin, indicating the significant benefit of incorporating KGs into LLM reasoning. In the LLM + KG category, FiDeLiS stands out as the best-performing method across all datasets, particularly when paired with GPT-4-turbo. For example, on WebQSP, FiDeLiS achieves 84.39% Hits@1 and 78.32% F1, surpassing ToG (81.84% Hits@1, 75.97% F1) and RoG (83.15% Hits@1, 69.81% F1). This improvement is consistent across other datasets, and even compared with some finetuning methods like DeCAF and RoG, FiDeLiS as a training-free method still demonstrate better performance. The consistent performance of FiDeLiS highlights its effective use of both the KG and LLM, as well as its optimization of hyper-parameters like beam width and depth. Overall, the results illustrate that FiDeLiS, leveraging advanced LLMs like GPT-4-turbo and KG-based reasoning, sets a new standard for performance in KG-related tasks.
@@ 4.1 Main Results
Hits@1 (%) Hits@1 (%) Acc (%)
No ablation FiDeLiS 79.32 63.12 67.34
[…]
w/o Path-RAG using vanilla retriever 72.35 57.11 59.78
using ToG 75.11 59.47 63.47
[…]
w/o DVBS w/o last step reasoning 75.68 59.45 63.72
w/o planning 76.23 60.14 64.13
w/o beam-search 60.35 49.78 61.87
w/o deductive-verifier 74.13 57.23 63.89
Table 2: Ablation Studies of FiDeLiS using model gpt-3.5-turbo-0125. $\Delta$ refers to the performance gap between each component and the entire method.
4.2 Ablation Study
Table 2 demonstrates the ablation study of FiDeLiS using the gpt-3.5-turbo-0125 model, highlighting the contributions of individual components (Path-RAG and DVBS) to overall performance. We conduct the ablation of the Path-RAG by replacing it with either a vanilla retriever or ToG Sun et al. (2023) as retriever. We find that using ToG shows slight improvements over the vanilla retriever but remains below using Path-RAG. Ablating DVBS components also leads to performance declines, particularly when beam search is removed, causing Hits@1 on WebQSP to drop sharply to 60.35%. The deductive verifier and last-step reasoning show moderate but noticeable impacts on performance. The effects are less pronounced on CR-LT, suggesting it is more tolerant of simpler methods. Overall, the results confirm the critical roles of Path-RAG and DVBS, especially beam search, in ensuring robust and accurate reasoning across domains.
@@ 4.2 Ablation Study
Vanilla Retriever w/ BM25 58.31 48.39 50.73
w/ SentenceBert 62.74 50.14 51.80
w/ E5 68.42 52.84 54.31
w/ Openai-Emb 72.35 57.11 59.78
[…]
Path-RAG w/ BM25 70.34 56.11 58.77
w/ SentenceBert 73.45 58.41 60.45
w/ E5 77.93 62.74 65.23
w/ Openai-Emb 79.32 63.12 67.34
Table 3: Performance of FiDeLiS with various embedding methods. * refers to text-embedding-3-small from OpenAI. We detail the tested embedding methods in §A.3.
4.3 Robustness Analysis
Robustness of Path-RAG.
Table 3 presents the performance of FiDeLiS compared to a vanilla retriever with different embedding methods. The results consistently show that FiDeLiS outperforms the vanilla retriever irrespective of the underlying embedding strategy. For instance, with Openai-Emb, the vanilla retriever achieves 72.35% on WebQSP, whereas Path-RAG reaches 79.32%, indicating a notable improvement. Similar performance gains are observed with the other embeddings. These improvements suggest that integrating graph connections can enhance retrieval effectiveness by providing more informative and contextually relevant information, thereby bolstering the overall robustness and accuracy of the method.
Effectiveness of Path-RAG.
We verify the effectiveness of the retrieval module Path-RAG with two baselines: (1) a vanilla retriever and (2) KAPING Baek et al. (2023) method. The vanilla retriever concatenates each entity with its relation to form a reasoning step and selects candidates based on cosine similarity with the query embeddings. In contrast, KAPING Baek et al. (2023) converts each triple into text and retrieves the top-$K$ similar triples based on semantic similarity. We quantify the retrieval performance using the coverage ratio (CR), defined as the percentage of the ground-truth reasoning steps being retrieved throughout the reasoning path extension (i.e., $\mathrm{CR}=\frac{N_{\mathrm{retrieved}}\cap N_{\mathrm{ground-truth}}}{N_{\mathrm{ground-truth}}}$). Table 4 illustrate the experimental setup and corresponding results. We find that compared with the baselines, our Path-RAG achieves a higher CR value and aligns better with the ground-truth paths. It demonstrates superior ability to capture connections that simpler retrieval models may overlook. This advantage is critical for guiding subsequent LLM processing toward relevant information, ultimately yielding more accurate and coherent answers.
Method Depth = 1 Depth = 2 Depth $>$ 3
Vanilla Retriever 59.34 52.17 47.31
KAPING Baek et al. (2023) 65.72 60.41 53.11
Path-RAG w/ keywords 72.61 69.38 62.78
Path-RAG w/o keywords 68.78 ($\downarrow$ 3.83) 65.27 ($\downarrow$ 4.11) 57.13 ($\downarrow$ 5.65)
Table 4: Analysis of the CR of reasoning paths over CWQ.
Methods WebQSP (hits@1) CWQ (hits@1)
Deductive Verification 79.32 63.12
Adequacy Verification (used in ToG) 74.13 57.23
Logit-based Scoring 73.47 54.78
Table 5: Analysis of different verification methods.
Path Error Analysis.
@@ 4.3 Robustness Analysis
To verify the faithfulness of our step-wise method, we conduct an error analysis regarding the whole reasoning path generation using RoG Luo et al. (2024a). We quantify the validity of reasoning path using validity ratio (VR), which is defined as the ratio of reasoning steps that existed in the KG to the total number of the reasoning steps in the output reasoning path (i.e., $\mathrm{VR}=\frac{N_{\mathrm{valid-steps}}}{N_{\text{all-steps}}}$). As shown in Figure 3, only 67% of generated reasoning steps are valid, while the remaining 33% of reasoning steps either have a format error or do not exist in the KG. This illustrates that the reasoning steps generated offer few guarantees about feasibility especially when multiple consecutive steps are combined into a reasoning path. While our method leverage step-wise verification to ensure that each of the reasoning step exist in the KG and logically connected.
Figure 3: Analysis of reasoning errors in RoG Luo et al. (2024a) over WebQSP.
@@
Table 6: Average depths of the generated reasoning paths. GT refers to ground-truth reasoning paths.
Effectiveness of Deductive-Verification.
To verify the effectiveness of deductive-verification mentioned in §3.2, we calculate the average depths of the generated reasoning paths as shown in Table 6. We find that by considering deductive verification, it consistently shows shorter and closer reasoning depths to ground-truth across all datasets compared to baseline. This implies that FiDeLiS may offer more precise termination signals and potentially more accurate reasoning paths. We also compare deductive-verification methods with other baselines in Table 5, like logit-based scoring that assign softmax probability scores to determine the endpoint of beam search process, and adequacy verification used in ToG Sun et al. (2023). Experiments show higher accuracy with deductive verification compared to adequacy verification and logit-based scoring, demonstrating its effectiveness in enhancing reasoning accuracy.
Dataset Method Hits@1 Runtime Token ##\##
@@
Table 7: Runtime efficiency of FiDeLiS per question.
[…]
Table 8: Case study of FiDeLiS. We highlight the wrong answers with red color, and correct answers with blue color.
This finding is further supported by a case study regarding a complex question of Iran’s government system, which blends elements of religion and democracy as shown in Table 8. While baseline methods such as GPT + CoT and RoG predominantly identified Iran as an “Islamic Republic” and ToG produce mixed responses, our approach—enhanced by deductive verification—delivers a reasoning path that is both concise and context-aware. The proposed verification mechanism not only streamlines the reasoning process but also ensures comprehensive coverage of grounded answers, demonstrating FiDeLiS’s strength in handling intricate questions.
4.4 Efficiency Analysis
To investigate the runtime efficiency and cost efficiency of FiDeLiS, we present a comparison regarding the average runtime, average token usage, average times of LLM calling per question in Table 7. We find that (1) our method shows superior efficiency compared to the ToG (which is also training-free), by reducing approximately 1.7x runtime costs. (2) Path-Rag component is critical in enhancing both the accuracy and efficiency of the model. Its ability to constrain potential path candidates effectively reduces unnecessary computational overhead, leading to quicker and more accurate results. To address concern regarding our method’s potential application in real-time scenarios, we also test our method using faster and more advanced LLMs. Table 7 shows that our method could be further accelerated with newer, faster models like GPT-4o or GPT-4-mini. The potential of the ongoing advancements in LLMs are expected to further enhance the scalability and efficiency of FiDeLiS, making it a practical development in challenging environments. More detailed analysis of bottleneck of computation of FiDeLiS can be further found in Appendix B.
@@ 5 Related Work
LLM Reasoning & Role of KGs.
Large language models (LLMs) demonstrate impressive capabilities in reasoning tasks but often generate hallucinated or factually incorrect outputs, particularly in complex, multi-step scenarios Huang et al. (2025); Li et al. (2024). This unreliability reduces their effectiveness in knowledge-intensive applications. Knowledge graphs (KGs) have emerged as a solution by offering structured, verifiable data that supports transparent and multi-hop reasoning Sui et al. (2024). Unlike document-based retrieval-augmented generation approaches, KGs provide direct access to relational facts, enhancing both interpretability and traceability Chen et al. (2024).
KG-enhanced LLM Reasoning.
KG-enhanced reasoning methods are generally categorized into retrieval-based and agent-based models. Retrieval-based approaches, such as DeCAF Yu et al. (2023), rely on text-based retrieval to select relevant information from KGs and jointly generate answers and logical forms, but their performance can degrade without precise retrieval mechanisms. In contrast, agent-based models, like ToG Sun et al. (2023), iteratively explore reasoning paths but suffer from high computational overhead. To address these limitations, recent methods like RoG Luo et al. (2024a) and GCR Luo et al. (2024b) have sought to integrate KG structure into LLM training or decoding to improve reasoning fidelity and explanation generation. To improve the faithfulness of the LLM reasoning, KD-CoT Wang et al. (2023) verifies sub-reasoning steps through external KGs to prevent errors during inference, while NSM He et al. (2021) employs a teacher-student architecture to learn intermediate supervision signals that guide reasoning.